### v1

# Investigation of hydrogen under-expanded jets in gaseous propulsion systems

Abstract
Underexpanded jets are present in various engineering applications; in recent years, they have gained special attention because of the development of gaseous-fueled propulsion systems. In these apparatuses, the direct injection of fuels such as hydrogen in innovative low-emission engines’ chambers induces turbulent under-expanded jets. In this study, we performed high-fidelity Large Eddy Simulations of under-expanded hydrogen jets to investigate these flows and provide valuable insights for developing injectors suitable for hydrogen and, more generally, gaseous-fueled propulsion systems. We initially assessed the method’s accuracy, evaluating the convergence and uncertainty of the numerical results and validating them against experimental particle image velocimetry and Schlieren data. The simulated jets, the Mach disc dimensions, and the resulting velocity field align closely with the experimental observations. Then we analyzed the jet structure for pressure ratios of 4 to 25 and examined the effects of the geometrical configuration of the nozzle on the characteristics of the air-fuel mixture obtained. We compared the jets resulting from a round-hole nozzle with annular ones resembling outward-opening injectors.
1 Introduction
Under-expanded jets are complex high-speed flows in multiple engineering applications, such as exhaust plumes from aircraft, rockets, missiles, and combustion chambers. Additionally, these jets can occur in natural phenomena, like volcanic eruptions (Orescanin et al., 2010; von der Linden et al., 2021; Carcano et al., 2013; FOX, 1974). Consequently, under-expanded jets have been extensively studied, particularly in aerospace applications. However, in the past decade, research has focused more and more on their role in the injection process of advanced propulsion systems, positioning this as an emerging field within fluid dynamics and engine simulations (Duronio et al., 2023, 2021). Therefore, it is crucial to develop a comprehensive understanding of these processes for propulsion systems (Onorati et al., 2022).
Modern propulsion systems often adopt gases such as hydrogen, propane, ammonia, and methane rather than liquids. Due to the significant pressure difference between the injector rail and the injection environment, supersonic conditions are achieved (Duronio et al., 2023; Vuorinen et al., 2014; Hamzehloo & Aleiferis, 2014b, 2016b), with the formation of underexpanded jets downstream of the injector nozzles, resulting in a distinctive flow field structure (Allocca et al., 2020; Samsam-Khayani et al., 2022).
Among the different combustibles, hydrogen is the subject of the most attention and research efforts; in this context, direct injection into the combustion chamber is one of the most promising technologies. Of course, the injection process plays a relevant role in the chain of events that take place in a propulsion system. Consequently, we must investigate the injectors and the resulting under-expanded jets to obtain the desired air-fuel mixture, efficient combustion, and reduced tailpipe emissions.
In recent years, researchers have focused their investigation on methane jets (Yosri et al., 2020; Banholzer et al., 2019; Bartolucci et al., 2020; Dong et al., 2018); comparatively less attention has been paid to hydrogen fuel, despite its growing scientific relevance. Experimental investigation exploits Schlieren imaging to capture the evolution of transient under-expanded hydrogen jets from gaseous fuel injectors, evaluating factors such as pressure ratio (PR), nozzle characteristics, and jet tip penetration (Coratella et al., 2024; Yeganeh et al., 2023a; Lee et al., 2021; Yeganeh et al., 2023b, 2022). Unlike jets of other species, examined using density maps and planar laser-induced fluorescence (PLIF), providing local values of fuel concentration and gas density (Yu et al., 2013b, 2011; Sakellarakis et al., 2021; Ni et al., 2022), experimental studies on hydrogen jets are often limited to qualitative observations, although detailed investigations are crucial for robust validation of new methods and CFD simulations of under-expanded jets.
Considering the intrinsic characteristics of the physical process under investigation, simulation tools play a crucial role in gaining deeper insights into the physics of under-expanded jets and supporting the development of simplified models for designing and optimizing injection devices. Common methodologies use both explicit (Onorati et al., 2022; Buttay et al., 2016; Hamzehloo & Aleiferis, 2019; Duronio et al., 2022) and implicit algorithms (Yosri et al., 2020; Rahantamialisoa et al., 2022; A & PG, 2014). Among these, the use of Large Eddy Simulation (LES) and high-order integration schemes with low numerical dissipation has proven to be particularly effective in replicating the characteristics of under-expanded jets and mixture formation (Hamzehloo & Aleiferis, 2014a; Vuorinen et al., 2014, 2013).
The literature covers extensive studies on under-expanded jets for gases such as methane, nitrogen, and air (Yu et al., 2013a; Banholzer et al., 2017; Traxinger et al., 2018; Banholzer et al., 2019; Xiao et al., 2019), with increasing recent interest in hydrogen (Anaclerio et al., 2022, 2023; Ballatore & van Oijen, 2024; Su et al., 2020). Hamzehloo & Aleiferis (2014b, 2016b, 2016a) explored near-nozzle characteristics of various under-expanded jets using a CFD code based on the Advection Upstream Splitting Method (AUSM) discretization for compressible flows, focusing on tip jet penetration, shear layers, and Mach disc structures. In their study, Ren & Wen (2020); Hecht & Panda (2019) examined cryogenic hydrogen () jets using both experimental and numerical methods to assess how the nozzle diameter and pressure ratio affect the structures of jet expansion. The authors utilized a two-dimensional (2D) model with weighted essentially non-oscillatory (WENO) schemes to enhance computational efficiency. A key distinction from typical setups is their use of a total pressure boundary condition instead of a high-pressure reservoir, a method also employed by (Zhang et al., 2019a). Further research has examined how real-fluid properties influence jet behaviour. Studies comparing the Redlich-Kwong and Peng-Robinson equations of state revealed that, under certain injection conditions, results differ significantly from those based on the ideal-gas law (Jin & Yao, 2021; Rahantamialisoa et al., 2023).
However, the most significant limitation of all these studies is the computational resources required. Historically, researchers studied the injection processes relying on Eulerian-Lagrangian CFD codes since they involved liquid fuels. The minimum grid sizing for these investigations is approximately equal to the nozzle diameter (Duronio et al., 2025a, b). Conversely, with gaseous fuels, a correct grid sizing, able to correctly represent the under-expanded jets, is of the order of magnitude of (Hamzehloo & Aleiferis, 2014a). Such a strict requirement implies a considerable increase in computational resources and severely limits the simulation’s feasibility. Careful meshing strategies, like multiple refinement regions and grid size gradually increasing downstream the nozzle exit section (Duronio et al., 2022; Hamzehloo & Aleiferis, 2016a), can help limit the computational load and allow simulations in acceptable time; nevertheless, increasing the grid dimension degrades the quality of the results, not correctly predicting the air/fuel mixture characteristics, especially when using RANS turbulence models (Wang et al., 2024).
Other studies simplify even more the problem by performing 2D simulations (Anaclerio et al., 2022; Ren & Wen, 2020; Zhang et al., 2019a). All these approaches are unreliable for developing new propulsion systems.
The paper reports a CFD investigation of hydrogen under-expanded jets for propulsion applications by high-fidelity GPU-accelerated simulations, adopting high-resolution grid sizing for the jet volume. To this purpose, we adopted AmRex PeleC (Henry de Frahan et al., 2024), a high-performance CFD solver for compressible reacting flows that, among the many characteristics, features Adaptive Mesh Refinement (AMR) and high-order order schemes.
We initially validated the simulations by comparison with measurements obtained by Particle Image Velocimetry (PIV) and Schlieren images of under-expanded jets (Ruggles & Ekoto, 2012; Edgington-Mitchell et al., 2014), then we investigate different injection configurations.
We studied different injection pressure ratios, evaluating the difference in the jet’s structure when hydrogen is adopted instead of air and the mixture formation process. We also focus on the nozzle characteristics, showing how the flow drastically changes with annular nozzles and hollow cone jets and how the mixing process is enhanced.
2 Mathematical Models and Numerical Method
2.1 Mathematical Models
The numerical simulation of the compressible multi-specie flow is performed by the integration of the following governing equations:
Here , and are the density, velocity, temperature, and pressure, respectively. is the total energy with representing the internal energy and is the mass fraction of the species. The viscous stress tensor, , under the Newtonian assumption, is given by:
where and are viscosity and the Lamè constant, respectively. is heat flux:
where is the thermal conductivity. is the diffusive transport flux of the k-th species:
where is the binary diffusivity coefficient between the species.
In all the simulations shown in the paper, we adopted the Dynamic Smagorinsky LES model (Germano et al., 1991; Smagorinsky, 1963) to accurately resolve the turbulent structures of the jet. For the sake of brevity, we are not reporting full details; the reader is referred to the original papers for a complete description. We supposed Schmidt and Prandtl numbers equal to 0.7 following Vuorinen et al. (2014); Hamzehloo & Aleiferis (2019); Duronio et al. (2022).
2.2 Numerical Method
The numerical integration of the governing equation (1) is performed using the solver AmRex PeleC (Henry de Frahan et al., 2024), which supports block-structured adaptive mesh refinement (AMR) and GPU parallelization (Zhang et al., 2019c). In the code, the inviscid fluxes in (1) are discretized using the unsplit piecewise parabolic method (PPM) with hybrid PPM WENO variants (Colella & Woodward, 1984; Motheau & Wakefield, 2020). The WENO reconstruction is performed with the 7th-order WENO-Z scheme of Balsara & Shu (2000). The diffusion fluxes are discretized in space with second-order centred differences. Temporal integration relies on a standard predictor-corrector approach (Henry de Frahan et al., 2024, 2022). Transport coefficients are evaluated from the CHEMKIN transport library and depend on temperature (Kee et al., 1986; Ern & Giovangigli, 1995). Thermodynamic properties are evaluated by means of the NASA7 polynomial parametrization (McBride, 2002).
3 Test cases: Characteristics and Numerical Setup
We investigated seven different case studies covering different morphologies of the under-expanded jets. Table 1 reports their main characteristics.
| Case N∘ | Nozzle | Gas/amb | [bar] | [bar] | [K] | Jet cone angle [∘] | ||
| Validation 1 | Round | H2/Air | 10 | 9.976 | 0.9976 | 292 | 1.5 | N/A |
| Validation 2 | Round | Air/Air | 4.2 | 4.25 | 1 | 288 | 15 | N/A |
| R1 | Round | H2/Air | 4.2 | 4.25 | 1 | 288 | 15 | N/A |
| R2 | Round | H2/Air | 12.6 | 12.8 | 1 | 288 | 15 | N/A |
| R3 | Round | H2/Air | 25.2 | 25.6 | 1 | 288 | 15 | N/A |
| A4 | Annular | H2/Air | 12.6 | 12.8 | 1 | 288 | 15 | 90 |
| A5 | Annular | H2/Air | 12.6 | 12.8 | 1 | 288 | 15 | 135 |
For verification and validation of the simulations, we considered the jets experimentally investigated by Ruggles & Ekoto (2012) and Edgington-Mitchell et al. (2014). Both are round jets from circular nozzles. The diameter is equal to for nozzle 1 and for nozzle 2. After these first verification cases, we performed other simulations in which we replicated case 2 by replacing air with hydrogen and then varied the pressure ratio from 4.2 to 25. The last two test cases A4 and A5 deal with annular nozzles, as shown in Figure 1(a). We defined them by maintaining the same inner diameter of cases R1-R3 and computing the outer diameter by imposing the same nozzle exit area. These configurations represent hollow cone injectors where the velocity of the gaseous fuel exhibits a radial component (), drastically changing the mixing process with the surrounding air. We studied two different values of the cone angle, representative of real prototypal devices (Montanaro, Alessandro et al., 2024).
We implemented the discrete equations on the 3D computational domain reported in figure 1(b).
The injection environment is a cubic box with sides equal to . Table 1 summarizes inflow and initial conditions. Validation cases 1 and 2 replicate the test conditions of Ruggles & Ekoto (2012) and Edgington-Mitchell et al. (2014). We also imposed open conditions at the bottom boundary while non-reflected boundary conditions at the sides of the box. We refined the mesh using fixed region refinements and Adaptive Mesh Refinement (AMR). In each region, the grid size obeys the law:
where is the dimension of the grid in the generic refinement region, is the base grid dimension, and is the refinement level. In all the simulations reported, we used and ; the size gradually decreases from the maximum to the minimum dimension (placed around the nozzle) with intermediate buffer levels. The region of maximum refinement around the nozzle has size .
We activated grid adaptation to guarantee the finest discretization of the whole jet and turned it on when one of the following conditions was verified:
1.
the velocity magnitude exceeds a threshold value , i.e.,;
2.
the maximum difference of the density in adjacent locations exceeds a threshold value equal to , i.e.
The integration interval depends on the jet’s time-lapse to reach the domain bottom, which is .
Finally, we built two coarser meshes for grid-dependence verification by doubling .
4 Verification and Validation
We assessed the convergence and uncertainty of the simulations using three grid levels, obtained as described at the end of section 3. We followed the procedure outlined in the classical paper by Roache (1997), in which the procedures now adopted by AIAA, ITTC, and IEEE are described and discussed in detail.
Table 2 reports the average values for the main fields evaluated along the jet’s axis and the relative grid convergence index (GCI) based on norms of the field variations, evaluated in the whole field at the end of the simulations.
| Case | 1 | Case | 2 | |||
| Average Value | GCI | Average Value | GCI | |||
| 0.25 | 1.54% | 0.47 | 1.47% | |||
| 0.19 | 1.27% | 0.43 | 1.27% | |||
| 1.10 | 2.52% | 1.17 | 2.84% |
To assess the adequacy of the chosen grid for Large Eddy Simulation, we evaluated the modelled kinetic energy and compared it to the total kinetic energy. This evaluation follows the methodology outlined in Di Mascio et al. (2022). Figure 2 illustrates the instantaneous ratio between the modelled and total kinetic energy computed once the jet reaches statistically steady conditions on the axial midplane.
The model kinetic energy is significantly lower than the resolved one, and the modeled-to-total kinetic energy ratio is less than 0.2 for most of the domain (the highest values appear in the shear layer). The ratio, therefore, is below the limit of 0.3, which guarantees adequate grid resolution for Large-Eddy Simulation (Pope & Pope, 2000).
Next, we compared the numerical results with experimental data. Figure 3 shows the time-averaged shape of the hydrogen jet and the Mach disc recorded using Schlieren imaging by Ruggles & Ekoto (2012) and the numerical results in terms of time-averaged density gradient for validation case 1.
As can be observed, the numerical simulation captures both the morphology of the hydrogen jet and the dimensions of the Mach disc. The jet has a peculiar conical shape bounded by a relevant mixing layer. Figure 4 compares the partial derivatives for the density and , relative to validation case 2.
Figure 5 reports the contours of mean axial and transverse flow velocities measured by Edgington-Mitchell et al. (2014) and computed in the present paper. We normalized all velocities with the value at the nozzle exit in the hypothesis of sonic conditions at the exit. The simulations correctly reproduce the morphology of the velocity field for both the axial and transverse components. Moreover, the computed Mach disc matches the PIV data, and the magnitude of the velocity is estimated correctly.
Figure 6 compares the time-averaged velocity profiles along the jet axis and several transverse sections. The agreement between the experimental data and the simulation is satisfactory.
5 Discussion of the results
We split the discussion into three parts. The first compares H2 and air under-expanded jets. Then, we investigate the pressure ratio’s effects on the hydrogen jets’ mixing process with the surrounding air. Finally, we examine the effects of the nozzle design (test cases A4 and A5) by varying the jet angle from 90∘ to 135∘.
5.1 Under-expanded H2 and Air jets
With this section, we aim to investigate the differences in an under-expanded jet’s flow structures (and, consequently, mixing characteristics) depending on the chemical species injected. Figure 7 reports the transient evolution of for Validation 2 and R1 cases. The conditions for the two test cases are identical, except that hydrogen is injected instead of air for test R1.
In both cases, we observe the classical structures of under-expanded jets, such as barrel shock, cap flow discontinuity, oblique shock waves, slipstream, reflected shocks, and so on (see, e.g., Duronio et al. (2022); Zhang et al. (2019b)). Nevertheless, we can underline several differences in the flow. First of all, the density of the H2 jet is lower than the ambient one. This will have a strong influence on the development of the jet. The initial vortex ring entraps much more surrounding fluid in the case of a hydrogen jet than with an air jet. A relevant amount of hydrogen moves aside on the exit section, producing remarkable vortical structures that create a hydrogen-enriched region in the radial direction (a sort of ”Coanda effect”). Figure 8 highlights this phenomenon by showing the hydrogen mass fraction in the near nozzle zone for two different time steps. This phenomenon is observed only with hydrogen and not air; experimental and CFD studies confirm this behaviour (Leick & Bartole, 2023; Merotto et al., 2024; Lejsek et al., 2024).
As the hydrogen jet evolves, we can observe the development and the evolution of a complex pattern of vortices (see figure 7 at and ), which is absent with air jets, with the presence of a strong shear layer. This complex mixing layer is essential in chemically reactive flow since it promotes interaction with the environmental gas.
We can explain the differences between the two jets by examining the vorticity equation
where we have the material of the vorticity on the left-hand side of 7, stretching and tilting in the first two terms on the right-hand side, the baroclinic term (the third term on the right), and the diffusion by viscosity in the last term. The baroclinic term produces vorticity when the density gradient is not aligned with the pressure gradient; this term is always present in compressible flows, but we expect strong effects when injecting light gas (hydrogen, in the present simulations) into heavy environment gas with respect to the situation of air-in-air or methane-in-air jets (Duronio & De Vita, 2024). Figure 9 reports the component of this term in non-dimensional form for the validation case 2 and R1 jets.
For the H2-in-air jet, the pressure gradient is almost parallel to the axis and directed to the left, while the density gradient is pointing outward with respect to the axis because its core is occupied by hydrogen, as shown by the arrows on the right of figure 9; therefore, the baroclinic term in the vorticity equation induces counterclockwise rotation in the upper part of the figure and clockwise rotation in the lower portion (as shown by the circular arrows in the picture) that push the hydrogen outward. On the contrary, in the air-in-air flow, the density gradient is pointing toward the jet axis, and consequently, the induced rotation is opposite to the one observed with the hydrogen flow. In addition, the magnitude of this term is much lower.
This baroclinic term justifies the completely different morphology of the outer jet between the single-species and the two-species flow. Nonetheless, the shock structure in the core is remarkably similar. Figure 10 is a zoomed view of the Mach disc for the two investigated cases performed after .
The figure shows that the structure of the primary shock cell is not affected by the species injected, as both the converging shocks and the normal shock have a similar morphology. As previously reported, the two-species jet exhibits a broader shear layer outside the jet core, where the hydrogen quickly mixes with the surrounding air. To quantify this aspect, we reported in figure 11 the magnitude of the vorticity vector after . The plot is a volumetric rendering where we removed a quarter of the jet to show its internal structure.
The vorticity is clearly higher for the hydrogen jet. In the near nozzle zone, both the jets present streamwise vortices outside the potential core with hairpin vortices downstream. As already underlined, these aspects are relevant in chemically reactive flows since the efficiency of the subsequent combustion process strongly depends on the mixing characteristics.
5.2 Effects of the pressure ratio on the H2 jets
Figure 12 reports the early stages of the jets R1, R2, and R3 evolution, where the pressure ratio equals 4.2, 12.6, and 25.2, respectively.
The early stages of the jets significantly differ when increasing the pressure ratio. The jet with the lower pressure ratio shows two converging shocks that merge at a point located on the jet’s axis and a pronounced primary vortex ring. When increasing the upstream pressure, the classical Mach disc shock appears, and the structure of the vortex rings changes. We can observe a primary vortex ring and a significant secondary vortex ring. A cap flow discontinuity delimits a distinguishable zone downstream of the Mach disc shock. The pressure ratio also has a significant effect on the shear layer. Indeed, when looking at the snapshots in figure 12, it can be concluded that a higher injection pressure promotes a broader shear layer, enhances the Coanda effect, and, as discussed below, helps form an ignitable air/fuel mixture. At t = (figure 13), the final under-expanded structures that characterized the jet can be observed.
Following the classification proposed by Franquet et al. (2015) and Duronio et al. (2023), the jet of case R1 is highly under-expanded, while the jets R2 and R3 are extremely (or very highly) under-expanded. Indeed, in the first case, the jet has a “barrel” structure that repeats several times downstream. The triple point can be recognized as the intersection of the intercepting shock, the Mach disc, and the reflected shock. Increasing the injection pressure, the structure of the jet is dominated by a single barrel shock. The Mach disc is no longer characterized by a normal shock, but a relevant curvature appears. The lack of a normal shock sequence promotes jet penetration and reduces the so-called potential core’s length, increasing the mixing between hydrogen and air; this aspect is shown in figure 14, where it can be observed that the zone of the jet characterized by the sequence of barrel shocks does not show any mixing activity, which is completely inhibited. The increase in the pressure ratio reduces the length of this zone.
Figure 15 shows a volumetric rendering of the vorticity magnitude for the three cases.
Vorticity is an important parameter in assessing the mixing activity: the higher the pressure ratio, the stronger the vorticity.
Obtaining a quantitative evaluation of the overall quality of the mixture is essential when dealing with combustion processes. Thus, we defined a discrete mass-weighted probability density function () as:
where:
N is the total cell count;
is the number of partitions; in this case K=100;
is the probability density function;
is the cell volume;
is the mixture density;
is the hydrogen mass fraction in the -th cell;
is the total injected mass at a certain instant;
Figure 16 reports the probability density function distribution and the mass injected for the three cases investigated at .
The PDFs for the three cases are very similar, showing a peak around ; of course, the amount of hydrogen mass injected grows with the injection pressure.
5.3 Effects of the nozzle shape
The nozzle design for gaseous injectors is essential to developing efficient injection devices. For this reason, we investigate two annular nozzles characterized by jet angles equal to 90∘ (case A4) and 135∘ (case A5). Figure 17 compares the jet’s transient evolution for the early stages of the simulated transient. As before, we plotted the .
The evolution is very different when compared with the round jet. At t =, the hollow-cone jets present a conical structure characterized by the angle imposed. The pressure inside the cone becomes gradually lower than the external one, and the jet continues its evolution by contracting towards the x-axis, as shown in figure 18 where we reported the time-averaged velocity vectors and the pressure field over on the axial plane.
Outside the core, in the near nozzle zone, air cannot enter within the jet due to the presence of a series of shock cells. On the other hand, we can observe a series of vortices rotating toward the central axis that drags the hydrogen into the jet’s core. This flow path extends for approximately 4D downstream of the nozzle exit section. The intensity of this phenomenon increases with the jet cone angle. As illustrated by figure 19, this circulation remains active until the jet spreads in the surroundings.
The subsequent injection stages are characterized by the jet advancing toward the right boundary of the domain with a velocity greater for jet A4 because of the smaller cone angle. Although the shape of the jet is similar, the cone angle modifies the hydrogen concentration, as shown in the figure by . Various experimental investigations confirm this evolution mechanism, and it is characteristic of hollow cone jets Lee et al. (2021); Coratella et al. (2024); Montanaro, Alessandro et al. (2024).
Figure 20 shows the Mach number for the two jets at t = .
The hollow conical structure of the jet determines a ring-shaped spatial arrangement of the shock waves. The single Mach disc observed with the circular nozzle disappears, but barrel shocks can also be recognized for test cases in A4 and A5; moreover, their size diminishes with the increasing jet angle. The number of these shock cells is larger for case A5. We can conclude that, when performing a hollow cone injection, the increase of the jet angle reduces the Mach number and the hydrogen velocity; therefore, with the same nozzle area, the injection of the same amount of requires a longer time.
To quantitatively assess the mixture quality obtainable with a hollow cone injection, in figure 21 we plotted the PDF and the average concentration into the domain as a function of time for the cases R2, A4, and A5. To have comparable results, we plotted the PDF at the time instant required to have an injected non-dimensional mass equal to 100 for all cases.
The improvement obtainable with the hollow-cone configurations, particularly with the one in test case A5, is remarkable. In the latter configuration, the high-concentration areas where the hydrogen is not flammable are almost entirely removed, and the average concentration decreases significantly. This is confirmed by figure 22, where we reported the hydrogen mass fraction at the time instant when the injected mass is the same.
The increased jet angle promotes the mixing with the surrounding air and almost completely removes the hydrogen-rich potential core. It can be seen that the region where the hydrogen mass fraction is close to one is shrunk, for case A5, to a hollow cone around the annular inflow section.
6 Conclusions
We investigated hydrogen under-expanded jets using Large Eddy Simulations of round and annular nozzles. We focused on the jet’s characteristics and, more importantly, the mixture formation process.
The simulations were verified and validated against experimental particle image velocimetry and Schlieren images when available. The simulated jets, the Mach disc dimensions, and the resulting velocity field closely align with experimental observations.
The analysis covers injection pressure ratios from 4.2 to 25.6 and examines the nozzle geometrical configuration’s effects on the air-fuel mixture’s characteristics. The main outcomes are:
The transient evolution of hydrogen under-expanded jets radically differs from the one observed for air. In particular, the mixing outside the jet core is enhanced by the baroclinic effect. On the contrary, the structure of the jet core is similar for the two flows in terms of shock structures.
The physical characteristics of the hydrogen trigger an intense interaction with the surrounding air. The hydrogen jet develops very high vorticity and engulfs a large air volume.
The pressure ratio is a key parameter in controlling the jet’s characteristics. It can be used to inject the proper amount of fuel in a reduced time.
Increasing the pressure ratio, the repeated shock cells disappear, and we have a unique Mach disc. Besides, the potential core becomes smaller for higher pressure ratios. As the potential core is the zone of the jet where there is no mixing activity when reducing it, the mixture quality improves.
Hollow cone jets have a drastically different morphology when compared to the classical round jets. Ring-shaped shock cells substitute the Mach disc, and the jet from highly under-expanded becomes weakly under-expanded. The Mach number decreases, and injecting the same amount of hydrogen requires a longer time.
The recirculation zone developing downstream the nozzle for approximately 4D significantly enhances the mixing activity, removes the hydrogen-rich core and gives a higher quality mixture.
A jet cone angle greater than 90∘ improves the mixture quality from the combustion point of view. With a jet cone angle equal to 135∘ we manage to quickly reduce the average hydrogen concentration already in the first phase of the injection and to remove the zones where the hydrogen concentration exceeds 0.4 completely.
Acknowledgment
This work has been funded by the European Union - NextGenerationEU, Mission 4, Component 1, under the Italian Ministry of University and Research (MUR) National Innovation Ecosystem grant ECS00000041 - VITALITY - CUP E13C22001060006.
Declaration of Interests
The authors report no conflict of interest.